How many electrons flow in 5 minute through an electric lamp which draw a current of 1 ampere given charge of an electron 1.6 * 10-19 coulomb

Grade:12th pass

2 Answers

Arun
4 years ago
When 1 Coulomb charge flow through a wire in 1 second then the current through the wire is 1 AMPERE.
 
I=Q/t
 
1 Ampere = 1 Coulomb /1 Second
Charge on 1 electron = 1.6 x 10^-19 Coulomb
By unitary method,
 
If 1.6 x 10^-19 Coulomb / Second (Ampere) = Current by 1 electron
then, 1( Coulomb / Second) or (Ampere) = 1 / (1.6 x 10^-19) electrons
i.e, 6.25 x 10^18 electrons.
 
Hencw for 5 minutes can also be calculated easily
Vikas TU
4 years ago
Dear student 
The above explanation is incorrect 
In 5 minutes there are 300 second 
So number of electrons will be 300 * 6.25 x 10^18  = 1875 * 10^18 electrons 
Good  Luck 
Cheers